40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jul 31 21:12:05 2013, manasa, Update, General, Pump down called off 
    Reply  Thu Aug 1 16:03:52 2013, Steve, Update, General, reopen BS & IOO IOOnorth.jpg
    Reply  Thu Aug 1 22:49:31 2013, manasa, Update, General, Green status after PR3 flipping investigated 
       Reply  Fri Aug 2 17:50:10 2013, Jenne, Update, General, PR3 wedge angle adjusted 
          Reply  Fri Aug 2 22:51:10 2013, Jenne, Update, General, All vent tasks complete, just need oplev check 
             Reply  Mon Aug 5 18:02:34 2013, manasa, Update, General, Close up list checked 
             Reply  Tue Aug 6 07:52:58 2013, manasa, Update, General, Ready to pump down MCspot_0806.pngOSEM_0806.pngOL_0806.pngQPD_0806.png
Message ID: 8958     Entry time: Thu Aug 1 22:49:31 2013     In reply to: 8948     Reply to this: 8960
Author: manasa 
Type: Update 
Category: General 
Subject: Green status after PR3 flipping investigated 

[EricQ, Koji, Manasa]

We opened the BS chamber to check the status of the green beams. The X green has 3 steering mirrors before they hit periscope1 and the Y green transmits through all the optics giving no way to steer it.

We agreed to start fixing the Y green. The wedge angle of PR3 is steering the transmitted beam away in both pitch and yaw. Since we are restricted only to yaw movement (done by moving the periscope), we want he wedge angle to be oriented in the yaw as well.

Right now, the wedge is oriented at about 20-30 deg off (The mark on the side of the mirror does not indicate the wedge). So we see a pitch as well as yaw misalignment in the transmitted beam. The pitch misalignment is making the beam fall off the mirrors in periscope2. 

We have decided to get the wedge angle set right for PR3 and redo the alignment for IR. Once we are aligned for the IR, we will modify the green layout.

ELOG V3.1.3-